# Bounds on the Privacy Amplification of Arbitrary Channels via the Contraction of $f_{\alpha}$-Divergence

### v3

@@ Abstract
We examine the privacy amplification of channels that do not necessarily satisfy any LDP guarantee by analyzing their contraction behavior in terms of $f_{\alpha}$-divergence, an $f$-divergence related to Rényi-divergence via a monotonic transformation. We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by Harremoës and Vajda [1]. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By applying these inequalities in a cross-channel setting, we arrive at strong data processing inequalities for $f_{\alpha}$-divergence that can be adapted to use-case specific restrictions of input distributions and channel. The application of these results to privacy amplification shows that even very sparse channels can lead to significant privacy amplification when used as a post-processing step after local differentially private mechanisms.
@@ I Introduction
An important property of local differential privacy (LDP) is its invariance to post-processing. Informally, the post-processing property states that whatever further processing is applied to the output of a system satisfying LDP will not weaken the overall system’s privacy guarantee. Mathematically, this property can be seen as a form of data processing inequality (DPI). As an extension of this observation, it is natural to ask about the privacy amplification of a post-processing operation. Basic results state that if a post-processing operation satisfying $\varepsilon_{2}$-LDP is applied to a system satisfying $\varepsilon_{1}$-LDP, the composition of both systems will satisfy at least $\min\{\varepsilon_{1},\varepsilon_{2}\}$-LDP. However, it is generally not know when arbitrary post-processing operations, that is, operations that themselves do not satisfy any LDP guarantee, may improve the overall systems LDP guarantee. We take steps to answer this question in this paper. In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [2, Chapter 33]). For pure LDP, the divergence measure of interest is the Rényi-divergence [3, 4] of order $\infty$, since
@@
Gilani et al. [5] suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$. With this framework at hand, studying the privacy amplification of a channel becomes equivalent to studying SDPIs for Rényi-divergences. While SDPIs have been extensively studied for $f$-divergences, results on Rényi-divergences are so far less explored. Although Rényi-divergence is not an $f$-divergence, it can be expressed as a monotonic transformation of a corresponding $f_{\alpha}$-divergence [6], a family of $f$-divergences parameterized by $\alpha$. This connection allows us to translate results on $f$-divergences to Rényi-divergences—results that are often easier to obtain.
@@ I-A Contributions
  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on privacy amplification in Section III.
@@
    • a simplification of the recently presented proof of Binette’s inequality [7] in [8] based on $E_{\gamma}$-divergence.
@@ I-B Related Work
The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [2, 9, 10, 11, 12, 13]. For Rényi-divergences, Polyanskiy and Wu [14] provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$. Recently, Jin et al. [15] show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid. In the context of privacy, Zamanlooy and Asoodeh [11] show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee. These bounds are further applied to mixing times in [16]. Related, it is shown in [17] that the LDP guarantee of a mechanism is closely related to its $f$-divergence contraction properties. The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [6] and subsequent works, including an applications to SDPIs in [18].
@@ II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [19] as
@@
The Rényi-divergence [3] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [2]:
@@
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [14, Section 3.2].
@@
For twice differentiable $f$, the relation shown in [20],
@@ II-B Strong Data Processing Inequalities
The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, in particular, it is shown in, e.g., [2, Theorem 33.6(c)] that
@@
It was shown by Ahlswede and Gács [13] that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$. Together with the data processing inequality for $f$-divergences, this implies that whenever (11) holds, we have $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. Next, it was shown in [15, Corollary 2] that
@@
Proposition 3 states that a channel does not contract the Rényi-divergence whenever there exist two input symbols that cause conditional distributions with disjoint support at the channel output. This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [21]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
@@
Heuristically, an edge $(x_{i},x_{j})$ between input symbols in the confusion graph indicates that there exists an output symbol $y\in\mathcal{Y}$ that can be caused by both $x_{i}$ or $x_{j}$ at the channel input. That is, a missing edge (compared to the complete version) indicates the existence of a pair of symbols that can be sent over the channel without being confused with each other. The confusion graph–along with its graph-theoretic dual–is extensively used in zero-error information theory, see e.g. [22, 23] for a detailed account of results and methods.
@@ IV Pinsker-type Inequalities for $f_{\alpha}$-divergence
[…]
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions. It was first shown by Binette [7] for general $f$-divergences. Here, we present a proof that utilizes the integral representation of $f$-divergences via the $E_{\gamma}$-divergence in (6). We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche and Tomamichel [8, Proposition 5.2]. However, the proof presented below slightly simplifies arguments, and we include it here as we believe it to be instructive.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [7, Theorem 1]).
@@
What remains is to bound $E_{\gamma}$ on the interval $I_{\geq 0}$, which yields an upper bound on $D_{f_{\alpha}}$ by (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [14]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $E_{\gamma}$ by the piecewise linear function
@@
The bound in [6, Theorem 35], [25, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
@@
For two mutually absolutely continuous probability distributions $P_{X}$ and $Q_{X}$ and any $\alpha>1$, we have
$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl{(}\text{TV}(P_{X}||Q_{X})\biggr{)}$ (38)
@@
For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [2]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$. Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
@@ V Strong Data Processing with Pinsker
[…]
@@
[…]
@@
Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$. ∎
This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ such that an operational bound on contraction is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. One such bound appears in the application of these results to privacy amplification, which we will present next.
@@ VI Privacy Amplification Bounds for RLDP
In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [27, 5]. Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [27] was recently suggested by Gilani et al. [5], where the authors refer to it as local Rényi differential privacy (LRDP). Note that the class of discrete mechanisms satisfying any (finite) RLDP guarantee for any $\alpha$ is equivalent to the class of discrete mechanisms satisfying some (finite) LDP guarantee. Specifically, this class only includes mechanisms with strictly positive entries in the mechanism matrix. i.e.,
@@
Further, assume that the channel $P_{Y|X}$ does not satisfy any finite RLDP guarantee. From the post-processing property of RLDP (which is just the DPI for Rényi-divergence), it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP channel $P_{X|W}$. Using our results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
@@
(a) $N=5$, $\varepsilon=\log 2$, $P_{Y|X}$
(b) $N=5$, $\varepsilon=\log 6$, $P_{Y|X}$
(c) $N=20$, $\varepsilon=\log 10$, $P_{Y|X}$
(d) $N=100$, $\varepsilon=\log 10$, $Q_{Y|X}$
@@
Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [28] satisfying $\varepsilon$-LDP, that is, let
@@
We examine the privacy amplification of the channels $P_{Y|X}$ and $Q_{Y|X}$ defined for some $N$ and $i,j\in[N]$ by
$P^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\nicefrac{{1}}{{2}},&\text{ if }i=j,\\ \nicefrac{{1}}{{2}},&\text{ if }i=j+1\mod N,\\ 0,&\text{otherwise}\end{cases}$ (59)
and for even values of $N$,
$Q^{(N)}_{Y|X=x_{i}}(y_{j})=\begin{cases}\frac{2}{N},&\text{if }1\leq i,j\leq\frac{N}{2}\\ \frac{2}{N},&\text{if }\frac{N}{2}<i,j\leq N\\ 0,&\text{otherwise}\end{cases}$ (60)
@@
In this setup, the successive application of both mechanisms to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the complete system compared to $P^{(N)}_{X|W}$ alone. We can numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the composition $P^{(N)}_{X|W}\circ P^{(N)}_{Y|X}$ or $P^{(N)}_{Y|X}\circ P_{X|W}$. As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight. Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $P^{(N)}_{Y|X}$,
$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{e^{\varepsilon}+1}{2}.$ (61)
On the other hand, given $Q^{(N)}_{Y|X}$, we have
$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=\frac{N+2e^{\varepsilon}-2}{N}.$ (62)
That is, the contraction behavior of the channels $P^{(N)}_{Y|X}$—and hence its privacy amplification—is independent of the alphabet size $N$. On the other hand, the amplification of the channels with $Q^{(N)}_{Y|X}$ crucially depends on the alphabet size. Interestingly, as long as $\varepsilon<\infty$, we see that asymptotically, $\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}(Y)=1$, that is, the privacy amplification of the channels $Q^{(N)}_{Y|X}$ becomes independent of the privacy guarantee of the system $P^{(N)}_{X|W}$. In particular, post-processing a randomized response mechanism with any finite LDP guarantee with a channel of a structure similar to $Q^{(N)}_{Y|X}$ will result in a system with perfect privacy as $N\to\infty$.
@@
Example 12 shows that even sparse channel matrices can lead to significant privacy amplification if applied as a post-processing step to an LDP system. Moreover, as long as the number of non-zero elements in the channel matrix grows linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows. While these insights are so far limited to randomized response mechanisms, they are interesting for two main reasons:
@@
First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [29, 30]. In this context, the results presented in this paper are encouraging. Specifically, if some form of LDP processing is done to the data, even a learning system without a formal LDP guarantee on its own might be able to significantly increase the privacy guarantee with respect to the data used to train it.
@@
Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [31]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.

### v4

@@ Abstract
We examine the privacy amplification of channels that do not necessarily satisfy any LDP guarantee by analyzing their contraction behavior in terms of $f_{\alpha}$-divergence, an $f$-divergence related to Rényi-divergence via a monotonic transformation. We present bounds on contraction for restricted sets of prior distributions via $f$-divergence inequalities and present an improved Pinsker’s inequality for $f_{\alpha}$-divergence based on the joint range technique by 5773031. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By applying these inequalities in a cross-channel setting, we arrive at strong data processing inequalities for $f_{\alpha}$-divergence that can be adapted to use-case specific restrictions of input distributions and channel. The application of these results to privacy amplification shows that even very sparse channels can lead to significant privacy amplification when used as a post-processing step after local differentially private mechanisms.
@@ I Introduction
An important property of local differential privacy (LDP) is its invariance to post-processing. Informally, the post-processing property states that whatever further processing is applied to the output of a system satisfying LDP will not weaken the overall system’s privacy guarantee. Mathematically, this property can be seen as a form of data processing inequality (DPI). As an extension of this observation, it is natural to ask about the privacy amplification of a post-processing operation. Basic results state that if a post-processing operation satisfying $\varepsilon_{2}$-LDP is applied to a system satisfying $\varepsilon_{1}$-LDP, the composition of both systems will satisfy at least $\min\{\varepsilon_{1},\varepsilon_{2}\}$-LDP. However, it is generally not know when arbitrary post-processing operations, that is, operations that themselves do not satisfy any LDP guarantee, may improve the overall systems LDP guarantee. We take steps to answer this question in this paper. In the context of data processing inequalities, this relates to the investigation of strong data processing inequalities (SDPIs), or the contraction of a channel with respect to a suitable measure of divergence (see [Polyanskiy_Wu_2025, Chapter 33]). For pure LDP, the divergence measure of interest is the Rényi-divergence [renyi1961entropy, van2014renyi] of order $\infty$, since
@@
gilani2024unifying suggest the notion of $\alpha$-Rényi-LDP as a relaxation of pure LDP, which bounds the maximum Rényi-divergence between the induced output distributions for some arbitrary order $\alpha$. With this framework at hand, studying the privacy amplification of a channel becomes equivalent to studying SDPIs for Rényi-divergences. While SDPIs have been extensively studied for $f$-divergences, results on Rényi-divergences are so far less explored. Although Rényi-divergence is not an $f$-divergence, it can be expressed as a monotonic transformation of a corresponding $f_{\alpha}$-divergence [7552457], a family of $f$-divergences parameterized by $\alpha$. This connection allows us to translate results on $f$-divergences to Rényi-divergences—results that are often easier to obtain.
@@ I-A Contributions
  • We derive a general condition under which channels fail to contract Rényi-divergence, establishing structural limits on the contraction of a channel in Section III.
@@
    • a simplification of the recently presented proof of Binette’s inequality [8630660] in [Hirche:2023caq] based on $E_{\gamma}$-divergence.
@@ I-B Related Work
The contraction behavior of $f$-divergences has been extensively studied; see, e.g., [Polyanskiy_Wu_2025, raginsky2016strong, 10.1134/S0032946020020015, 10206578, anantharam2013maximal, ahlswede1976spreading]. For Rényi-divergences, polyanskiy2015dissipation provide upper bounds in the Gaussian setting using the Hockeystick-divergence $E_{\gamma}$. Recently, 10619367 show that the standard total-variation-based upper bound fails for Rényi-divergence, though the universal $\chi^{2}$-based lower bound remains valid. In the context of privacy, 10206578 show that if a mechanism satisfies local differential privacy, strong data processing inequalities can be derived in terms of a channel’s LDP guarantee. These bounds are further applied to mixing times in [zamanlooy2024mathrm]. Related, it is shown in [9517999] that the LDP guarantee of a mechanism is closely related to its $f$-divergence contraction properties. The $f_{\alpha}$-divergence framework enables the use of $f$-divergence inequalities in the Rényi setting, as developed in [7552457] and subsequent works, including an applications to SDPIs in [sason2019data]. The privacy amplification of various mechanisms in the central RDP setting is studied in the line of work [balle2018privacy, balle2019privacy, balle2020privacy].
@@ II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [csiszar1967information] as
@@
The Rényi-divergence [renyi1961entropy] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [Polyanskiy_Wu_2025]:
@@
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [polyanskiy2015dissipation, Section 3.2].
@@
For twice differentiable $f$, the relation shown in [cohen1998comparisons],
@@ II-B Strong Data Processing Inequalities
The contraction coefficient of the $\chi^{2}$-divergence is closely related to $\rho_{m}$, and it is shown in, e.g., [Polyanskiy_Wu_2025, Thm. 33.6(c)] that
@@
It was shown by ahlswede1976spreading that $s(P_{Y|X})<1$ if and only if for every pair $x,x^{\prime}$, it holds that $\textnormal{supp}(P_{Y|X=x})\cap\textnormal{supp}(P_{Y|X=x^{\prime}})\neq\emptyset$. Together with the data processing inequality for $f$-divergences, this implies that whenever (11) holds, we have $\eta_{\chi^{2}}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. Next, it was shown in [10619367, Corollary 2] that
@@
Proposition 3 states that a channel does not contract the Rényi-divergence whenever there exist two input symbols that cause conditional distributions with disjoint support at the channel output. This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [diestel2017]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
@@
Heuristically, an edge $(x_{i},x_{j})$ between input symbols in the confusion graph indicates that there exists an output symbol $y\in\mathcal{Y}$ that can be caused by both $x_{i}$ or $x_{j}$ at the channel input. That is, a missing edge (compared to the complete version) indicates the existence of a pair of symbols that can be sent over the channel without being confused with each other. The confusion graph—along with its graph-theoretic dual—is extensively used in zero-error information theory, see e.g., [1056798, 720537] for a detailed account of results and methods.
@@ IV Pinsker-type Inequalities for $f_{\alpha}$-divergence
In this section, we present a generalization of the reverse Pinker’s inequality and Pinsker’s inequality for Rényi-divergences. These inequalities will then be used to derive “SDPI-style” bounds in Section V. In most of what follows, we will work with the $f_{\alpha}$-divergence as a substitute for Rényi-divergence. Due to the monotonic transformation in (4), conversion between the two can be done easily at any point.
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions. It was first shown by 8630660 for general $f$-divergences. Here, we present a proof that utilizes the integral representation of $f$-divergences via the $E_{\gamma}$-divergence in (6). We remark that the same technique was recently used to proof Binette’s inequality in a quantum context by Hirche:2023caq. However, the proof presented below slightly simplifies arguments, and we include it here as we believe it to be instructive.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [8630660, Theorem 1]).
@@
What remains is to bound $E_{\gamma}$ on the interval $I_{\geq 0}$, which yields an upper bound on $D_{f_{\alpha}}$ by (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [polyanskiy2015dissipation]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $E_{\gamma}$ by the piecewise linear function
@@
The bound in [7552457, Theorem 35], [7360766, Theorem 3] is obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}(X)=0$.
@@
For any $\alpha>1$ and any two mutually absolutely continuous probability measures $P_{X}\ll\gg Q_{X}$, we have
$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl(\text{TV}(P_{X}||Q_{X})\biggr),$ (38)
@@
For any $\alpha>1$, the bound in Theorem 8 is attained by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [Polyanskiy_Wu_2025]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ holds for all $t>0$. Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
@@ V Strong Data Processing with Pinsker
The above results can now be used to reason about the data processing properties of channels by considering the presented inequalities “across” a given channel. That is, we can modify the bound presented in Theorem 6 to yield an upper bound on $D_{f_{\alpha}}(P_{Y}||Q_{Y})$ in terms of $\text{TV}(P_{X}||Q_{X})$. To this end, consider
@@
as the “cross-channel” modification of the crucial extremal quantities in Theorem 6. We get the following result.
@@
Follows from Theorem 6 with $P_{Y}=P_{Y|X}\circ P_{X}$ and $Q_{Y}=P_{Y|X}\circ Q_{X}$, the inequality in Theorem 8, the inverse given in Remark 10 and since by definition, $\text{TV}(P_{Y}||Q_{Y})\leq\eta_{TV}(P_{Y|X},\mathcal{P})\text{TV}(P_{X}||Q_{X})$ if $P_{X},Q_{X}\in\mathcal{P}$. ∎
This result stresses the relation between $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ as formulated in (42) and (43) and the contraction behavior of a channel in terms of Rényi-divergence: If $\Gamma_{\max}(Y)-\Gamma_{\min}(Y)$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}(Y)$ and $\Gamma_{\min}(Y)$ so that an operational bound on this contraction is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. One such bound can be obtained in the application of the above results to privacy amplification, which we will present in the following section.
@@ VI Privacy Amplification Bounds for RLDP
In this section, we show how the above technical tools can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [mironov2017renyi, gilani2024unifying]. Consider the Markov chain WXYW-X-Y, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a mechanism satisfying $(\varepsilon,\alpha)$-RLDP, [^1] The notion of Rényi local differential privacy as a local adaption of Rényi differential privacy [mironov2017renyi] was recently suggested by gilani2024unifying, where the authors refer to it as local Rényi differential privacy (LRDP). Note that the class of discrete mechanisms satisfying any (finite) RLDP guarantee for any $\alpha$ is equivalent to the class of discrete mechanisms satisfying some (finite) LDP guarantee. Specifically, this class only includes mechanisms with strictly positive entries in the mechanism matrix. i.e.,
@@
Further, assume that the channel $P_{Y|X}$ does not satisfy any finite RLDP guarantee. From the post-processing property of RLDP (which is just the DPI for Rényi-divergence), it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there are certain $P_{Y|X}$ with zero-valued elements that improve the privacy guarantee of $P_{Y|W}$ when cascaded with the RLDP mechanism $P_{X|W}$. Using the results above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
@@
(a) $N=5$, $\varepsilon=\log 2$, $k=2$
(b) $N=5$, $\varepsilon=\log 6$, $k=2$
(c) $N=20$, $\varepsilon=\log 10$, $k=2$
(d) $N=100$, $\varepsilon=\log 10$, $k=50$
@@
Let $P^{(N)}_{X|W}$ be the $N\times N$ randomized response mechanism in [extremalmechanismLong] satisfying $\varepsilon$-LDP, that is, let
@@
For some $m,k>0$, let $N=mk$. We examine the privacy amplification of the channels $Q^{(m,k)}_{Y|X}$ defined by
$Q_{Y|X}^{(m,k)}\coloneqq\text{diag}(\underbrace{U^{(k)},\dots,U^{(k)}}_{m\text{ times}}),$ (59)
where $U^{(k)}=\frac{1}{k}\mathbbm{1}_{k}\mathbbm{1}_{k}^{T}$ is a “uniform” block of size $k\times k$. In this setup, the successive application of $P_{X|W}^{(N)}$ and $Q_{Y|X}^{(m,k)}$ to the private data $W$ will increase randomization, and hence it is reasonable to expect an improved privacy guarantee of the composed system compared to $P^{(N)}_{X|W}$ alone. Computing the values of $\Gamma_{\max}$ and $\Gamma_{\min}$ in this setup, we find that given the channel $Q^{(m,k)}_{Y|X}$, we have
$\Gamma_{\max}(Y)=\Gamma_{\min}(Y)^{-1}=1+\frac{e^{\varepsilon}-1}{k}=1+\frac{m(e^{\varepsilon}-1)}{N}.$ (60)
Interestingly, we observe that for any $\varepsilon>0$, and any fixed ratio of non-zero elements $\nicefrac{{k}}{{N}}=\nicefrac{{1}}{{m}}$, we have
$\lim_{N\to\infty}\Gamma_{\max}(Y)=\lim_{N\to\infty}\Gamma_{\min}=1,$ (61)
which implies that for any $\alpha>1$, we have
$\varepsilon^{*}=\text{RLDP}^{(\alpha)}\Big(Q^{(\nicefrac{{N}}{{k}},k)}_{Y|X}\circ P_{X|W}^{(N)}\Big)\longrightarrow 0.$ (62)
@@
That is, post-processing a randomized response mechanism of any finite LDP guarantee with a channel constructed in the above way will result in a system with perfect privacy as $N\to\infty$ whenever the ratio of non-zero elements remains fixed as $N$ grows. We can further numerically determine the bound in (53) assuming the upper bound $\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. The resulting RLDP values are shown in Figure 2 alongside the RLDP guarantee of $P^{(N)}_{X|W}$ (the best bound previously known) and the true RLDP guarantee of the compositions for different choices of $N$ and $k$. As expected, the bound tightens as $\alpha$ grows, since both of the presented Pinsker-type bounds are asymptotically tight in the parameter $\alpha$.
@@
Example 12 shows that even sparse channel matrices can lead to significant privacy amplification if applied as a post-processing step to an LDP system. Moreover, as long as the number of non-zero elements in the channel matrix grows at least linearly with $N$, the privacy guarantee of the system will approach prefect privacy as the alphabet size grows. While these insights are so far limited to randomized response mechanisms, they are interesting for two main reasons:
@@
First, the results motivate us to further investigate arbitrary systems that apply some form of randomization in terms of their effect on LDP guarantees; it has been previously observed that local differential privacy constraints on a system can be severely limiting, especially in large-scale learning applications [10.1145/3433638, 10.5555/3454287.3455674]. In this context, the results presented in this paper are encouraging. Specifically, if some form of LDP processing is done to the data, even a learning system without a formal LDP guarantee on its own might be able to significantly increase the privacy guarantee with respect to the data used to train it.
@@
Second, the results are promising for addressing the common issue of large values of $\varepsilon$ in $\varepsilon$-LDP guarantees in practical systems [dwork2019exposeyourepsilons]: Instead of fine tuning the practical application itself to improve its privacy guarantee (which might be intractable), practitioners might be able to apply simple post-processing steps to an application output in order to significantly improve the systems privacy guarantee, while retaining much of the original systems performance.